# GenM3: Generative Pretrained Multi-path Motion Model for Text Conditional Human Motion Generation

### v1

@@ 3.2 Training Strategy
where $\beta$ is the hyper-parameter to balance the losses. Additionally, we utilize moving averages for codebook updates and resets for inactive codes.
@@ 5.1 Datasets and Implementation
We collected and processed a total of 11 motion capture datasets [1, 42, 29, 21, 44, 27, 10, 3, 23, 20, 18], aligning each to the HumanML3D dataset’s format for consistency in training and evaluation. We use all the motion data for pretraining. For text-conditional training, we used datasets that include textual modality, specifically HumanML3D [10], HuMMan [3], HIMO [21], BABEL [27], and IMHD [44].
@@ 5.3 Ablation Study
We conducted experiments by pretraining MMM [26], T2M-GPT [38] and our baseline (refers to our method without the Multi-path, Multi Expert, and Motion Descriptor components) on our integrated dataset in Fig. 6. It can be seen that pretraining on our integrated dataset yields the greatest improvement (35.21%) for our method. This can be attributed to the multi-expert design in our MEVQ-VAE, which enables the model to adapt to the heterogeneity of various datasets. Moreover, the Motion Pathway in MMT leverages a multi-expert combination to simultaneously learn motion patterns from different datasets.
@@
Impact of Expert Count
Table 5: Comparisons of dense MoE and sparse MoE in the motion pathway.
[…]
Sparse MoE 0.058 0.799 $\boldsymbol{9.676}$
Dense MoE $\boldsymbol{0.046}$ $\boldsymbol{0.804}$ 9,675
Dense MoE v.s. Sparse MoE
@@ 6 Conclusion
In this supplementary material, we provide more information that could not be included in the main manuscript because of space limit. We first provide more details about our constructed dataset and implementations in Sec. 7 and 8. Sec. 9 shows the visualization results of generated motion corresponding to different text inputs. Finally, we discuss limitations of our proposed MMGPT in Sec. 10.
@@
BEATv2 [20] 9.43 87.29 -
CIRCLE [1] 1.07 9.91 -
EgoBody [42] 0.40 3.70 -
GRAB [29] 0.40 3.74 -
@@ 7 Details of the Dataset
We collected and processed 11 high-quality motion capture datasets, including CIRCLE [1], Egobody [42], GRAB [29], HIMO [21], IMHD [44], BABEL [27], HumanML3D [10], HuMMan [3], MPI-INF-3DHP [23], InterHuman [18] and BEATv2 [20]. All datasets were processed at 30 fps, with the duration of each sample ranging from 2 to 10 seconds. Tab.  6 illustrates the number of frames included in each dataset after processing.

### v2

@@ 3.2 Training Strategy
where $\beta$ is the hyper-parameter to balance the losses. Additionally, we utilize moving averages for codebook updates and resets for inactive codes [38].
@@ 5.1 Datasets and Implementation
We collected and processed a total of 11 motion capture datasets [1, 44, 30, 21, 46, 28, 10, 3, 23, 20, 18], aligning each to the HumanML3D dataset’s format for consistency in training and evaluation. We use all the motion data for pretraining. For text-conditional training, we used datasets that include text modality: HumanML3D [10], HuMMan [3], HIMO [21], BABEL [28], and IMHD [46].
@@ 5.2 Comparison
Table 3: Comparison of different VQ methods.
Method VQ MEVQVAE RVQ [12] MERVQ G-RVQ [37] FSQ [24]
FID 0.098 0.048 0.043 0.032 0.045 0.057
Comparison of VQ Method
Tab. 3 compares different quantization methods, including RVQ [12], G-RVQ [37] and FSQ [24]. Our multi-expert design adapts to diverse dataset distributions, substantially outperforming standard VQ. RVQ achieves the best scores but requires an extra Residual Transformer to model residuals [12]. Integrating our Multi-Expert into RVQ (MERVQ) further boosts performance, demonstrating its generalizability.
@@ 5.3 Ablation Study
We conducted experiments by pretraining MMM [27], T2M-GPT [40] and our baseline (refers to our method without the Multi-path, Multi Expert, and Motion Descriptor components) on our integrated dataset in Fig. 6. It can be seen that pretraining on our integrated dataset yields the greatest improvement (35.21%) for our method. This can be attributed to the multi-expert design in our MEVQ-VAE, which enables the model to adapt to the heterogeneity of various datasets. Moreover, the Motion Pathway in MMT leverages multiple experts to simultaneously learn motion patterns from different datasets.
@@
Impact of Expert Count
Table 5: Comparisons of dense MoE and sparse MoE.
[…]
Sparse MoE 0.058 0.799 $\boldsymbol{9.676}$
Dense MoE $\boldsymbol{0.046}$ $\boldsymbol{0.804}$ 9,675
Dense MoE v.s. Sparse MoE
@@ 6 Conclusion
7 Acknowledgment
This work is partially supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2025A03J3656).
@@ 7 Acknowledgment
In this supplementary material, we provide more information that could not be included in the main manuscript because of space limit. We first provide more details about our constructed dataset and implementations in Sec. 8 and 9. More results and ablation studies are conducted in Sec. 10 Sec. 11 shows the visualization results of generated motion corresponding to different text inputs. Finally, we discuss limitations of our proposed GenM3 in Sec. 12.
@@
BEATv2 [20] 9.43 87.29
CIRCLE [1] 1.07 9.91
EgoBody [44] 0.40 3.70
GRAB [30] 0.40 3.74
@@ 8 Details of the Dataset
We collected and processed 11 high-quality motion capture datasets, including CIRCLE [1], Egobody [44], GRAB [30], HIMO [21], IMHD [46], BABEL [28], HumanML3D [10], HuMMan [3], MPI-INF-3DHP [23], InterHuman [18] and BEATv2 [20]. All datasets were processed at 30 fps, with the duration of each sample ranging from 2 to 10 seconds. Tab. 7 illustrates the number of frames included in each dataset after processing.
@@ 9 Implementation Details
10 More Results
Table 8: Inference speed per sample (on a RTX4090 GPU) with different iteration settings during masked decoding.
Method FID Infer. time (ms) R-Precision Diversity
GenM3 (iter.=1) 2.242 23.73 0.679 8.477
GenM3 (iter.=5) 0.054 67.49 0.785 9.375
GenM3 (iter.=10) 0.046 113.22 0.804 9.675
GenM3 (iter.=15) 0.060 188.12 0.805 9.719
T2M-GPT 0.160 108.03 0.770 9.653
Inference Speed
During decoding, all tokens are decoded in parallel at each iteration. Tab. 8 shows how iteration count (10 for default settings) relates to inference time, and compares this with the AR model T2M-GPT [40].
Table 9: Results on HumanML3D under varying data proportions.
Data Proportion 0% 35% 70% 100%
FID 0.071 0.060 0.050 0.046
Influence of Dataset Size
We have evaluated our backbone pre-training on varying proportions of the dataset (HumanML3D dataset is fully used). We can conclude that the quality of the generation scales with the size of dataset (see Tab. 9). We believe that integrating additional motion data can unlock further model potential, which will be the focus of our future work.
Table 10: Results on HumanML3D (annotated by FineMotion).
Method FID R.Top1 R.Top2 R.Top3 MMDist. Diversity
MMM 24.7 0.060 0.113 0.159 5.85 6.01
GenM3 17.2 0.072 0.135 0.191 5.30 6.00
GenM3 12.9 0.087 0.154 0.218 5.01 6.25
Zero-shot Generation
We added comparisons on unseen HumanML3D annotations from FineMotion [^1] https://github.com/BizhuWu/FineMotion (see Tab. 10) to further prove GenM3’s generalization ability. Since the re-annotated descriptions differ significantly from those in the original HumanML3D, the zero-shot generation performance is generally not ideal. However, our method achieves better results compared to MMM [27]. By comparing the results of GenM3 and GenM3, it can be concluded that training on a larger dataset (with more text annotations and more motion data) improves the model’s generalization ability.